OpenStudy (anonymous):

Determine whether Rolle's Theorem can be applied to f on the indicated interval. If Rolle's Theorem can be applied, find all values of c in the interval. f(x)= x^2 - 2x -3/ (x +2) show all steps please

OpenStudy (anonymous):

What is the interval given?

OpenStudy (anonymous):

[-1,3]

OpenStudy (anonymous):

The only value that would exclude the possibility for Rolle's theorem on this problem ix x=-2 (the value that would produce a divide by zero). You can use Rolle's Theorem because your interval is from -1 to 3 and does not include x=-2.

OpenStudy (anonymous):

How do you solve this?

OpenStudy (anonymous):

Check that f(-1)=f(3).

OpenStudy (anonymous):

If it does, take the derivative of the function and set it equal to zero to find the x location where the function has a horizontal tangent... that is your c value.

OpenStudy (anonymous):

how would you take the derivative of x^2 - 2x -3/ (x + 2)?

OpenStudy (anonymous):

Quotient rule:\[\frac{d}{dx}[\frac{f(x)}{g(x)}]=\frac{f'(x)g(x)-f(x)g'(x)}{[g(x)]^{2}}\]

OpenStudy (anonymous):

show steps

OpenStudy (anonymous):

\[f(x)=x ^{2}-2x-3\]\[f'(x)=2x-2\]\[g(x)=x+2\]\[g'(x)=1\]Now place these into the formula for the quotient rule above.

OpenStudy (anonymous):

Let me know what you get.

OpenStudy (anonymous):

where are getting g(x)?

OpenStudy (anonymous):

The denominator.

OpenStudy (anonymous):

Sorry, I re-used f(x)... most people look at that as the numerator for a quotient rule and g(x) as the denominator. I missed that f(x) was already used for your initial function. The idea is that you have a numerator function (usually f) and a denominator (usually g).

OpenStudy (anonymous):

I hope that you have seen this before if you are working on Rolle's Theorem. Usually people do a whole chapter on derivative techniques before they tackle Rolle's.

OpenStudy (anonymous):

I can re-state:\[\frac{d}{dx}[\frac{n}{d}]=\frac{n'd-nd'}{d ^{2}}\]\[n=x ^{2}-2x-3\]\[n'=2x-2\]\[d=x+2\]\[d'=1\]\[\frac{d}{dx}[\frac{n}{d}]=\frac{(2x-2)(x+2)-(x ^{2}-2x-3)(1)}{(x+2)^{2}}\]Now simplify.